Problem 3. A free-standing ambulatory care center averages $70 in charges per patient. Variable costs are approximately $12 per patient, and fixed costs are about $1.5 million per year. Using this data, how many patients must be seen each day, assuming a 365-day operation, to reach the break-even point?
Answer #1

Total contribution per patient = Ambulatory expense- Variable costs

= 70-12 = 58

Break even point -Annual = Fixed cost/ Contribution per patient

= 1500000/58

= 25862.07

Daily Break even = 25862.07/365 = 70.85

rounded off to 71 patients